OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). Answer y = 2x - 7 y = -1/2 x - 2 y = 2x - 2 y = -1/2 x - 7

OpenStudy (angela210793):

parallel means tht they have the same slope=2 so the eq ur looking is y-y0=m(x-x0) y-3=2(x-5) y=2x-7

OpenStudy (anonymous):

there is no negative one sorry